1. Given the model, ( ) ( ) 0 0 – 0, 0 1 C a b Y T Y C I G T d tY a b = +   = =  + + + a) How many endogenous variables are there? b) Find the equilibrium values of Y T C


1
Expert's answer
2021-03-14T19:50:54-0400

a) Endogenous variables are C, a, b, T, I, G, Yd, t.

b) In equilibrium Y = C + I + G, C = C0 + c(Y - T), T = tY.